Find the general solution of
y" + 3y' + 2y = 1/(1+e^x)
1
Expert's answer
2020-07-23T18:13:14-0400

We solve a second order differential equation of the type:

ay''+by'+cy=f(x)

where a=1

b=3

c=2

f(x)=1/(1+e^x)

The solution will consist of common and particular solutions and have a type:

ygen=ycom+ypart

To find a common solution we solve second order differential equation y''+3y'+2y=0 and find roots of polynomial r2+3r+2=0

Its roots  

"r_1={{-3+\sqrt{3^2-4\sdot1\sdot2}}\over{2\sdot1}}={{{-3+\sqrt{1}}\over{2}}={{-3+1}\over{2}}}={-1}"

"r_2={{-3-\sqrt{3^2-4\sdot1\sdot2}}\over{2\sdot1}}={{{-3-\sqrt{1}}\over{2}}={{-3-1}\over{2}}}={-2}"

​Roots are complex therefore the common solution will be "y_{com}=C_1e^{-{x}}+C_2e^{-2{x}}"

We will use the method of variation of parameters to find a particular solution

"\begin{cases}
C'_1(x)y_1+C'_2(x)y_2=0 \\
C'_1(x)y'_1+C'_2(x)y'_2={f(x)\over {a_0}}
\end{cases}"

"y_1=e^{-{x}}\\
y_2=e^{-{2x}}\\
a_0=1"

finally we get

"\begin{cases}
C'_1(x)e^{-{x}}+C'_2(x)e^{-{2x}}=0 (1)\\
-C'_1(x)e^{-{x}}-2C'_2(x)e^{-{2x}}={1\over {1+e^{{x}}}} (2)
\end{cases}"

add (2) to (1) and we get

"-2C'_2(x)e^{-{2x}}={1\over {1+e^{x}}}\\
C'_2(x)=-{e^{2x}\over {1+e^{x}}}"

We will make some changes and simplifications

"C'_2(x)=-{e^{2x}+e^{x}-e^{x}\over {1+e^{x}}}=-{e^{x}(e^{x}+1)\over {e^{x}+1}}+{e^{x}\over {e^{x}+1}}=-e^{x}+{e^{x}\over {e^{x}+1}}"

Let us integrate C'2

"C_2(x)=\lmoustache{(-e^{x}+{e^{x}\over {e^{x}+1}})dx}=-e^{x}+ln\vert{{e^{x}+1}}\vert+C_2"

we multiply (1) by 2, add to (2) and get

"C'_1(x)e^{-{x}}={1\over {1+e^{x}}}\\
C'_1(x)={e^{x}\over {1+e^{x}}}"

Let us integrate C'1

"C_1(x)=\lmoustache{({e^{x}\over {e^{x}+1}})dx}=ln\vert{{e^{x}+1}}\vert+C_1"

Now we add these values in our general solution for y and will get

"y_{gen}=(-e^{x}+ln\vert{{e^{x}+1}}\vert+C_1)e^{-{x}}+e^{-{2x}}(ln\vert{{e^{x}+1}}\vert+C_2)"

Finally we get

"y_{gen}=1+e^{-{x}}ln\vert{{e^{x}+1}}\vert+C_1e^{-{x}}+e^{-{2x}}ln\vert{{e^{x}+1}}\vert+C_2e^{-{2x}}"